# Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach

### v1

@@ 6.3 Forecasting Results
Table 6: Forecast Inference Compared to Net-GO-GARCH: Bootstrap Confidence Intervals and Diebold and Mariano (2002) Tests (Sorted Worst to Best, Standard DCC-GARCH Excluded)
Model RMSFE CI MAFE CI DM $p$-value
Lower Upper Lower Upper
Std. GO-GARCH 5.7422 6.4045 5.3793 5.9524 $<0.001^{***}$
Piccolo-Diss 5.9414 6.6492 5.0637 5.6824 $0.0106^{**}$
Correlation-Diss 3.7825 4.7440 2.8536 3.5627 $<0.001^{***}$
Euclidean-Diss 3.5847 4.5018 2.7187 3.3796 $<0.001^{***}$
Std. CCC-GARCH 2.3063 2.9416 1.6726 2.1596 $0.0018^{***}$
Net. CCC-GARCH 2.1806 2.6795 1.6545 2.0391 $0.0811$
Net. DCC-GARCH 2.1806 2.6794 1.6547 2.0394 $0.1029$
Benchmark: Net-GO-GARCH 2.1797 2.6766 1.6541 2.0376
Notes: 95% bootstrap confidence intervals based on 2000 replications. Diebold–Mariano (DM) tests compare predictive accuracy against the benchmark network log-ARCH model with GO-GARCH-based distances. CW results were omitted to focus on non-nested comparisons. Significance levels: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$.
@@
Table 7: Forecast Comparison: Best Network Models vs. Standard DCC-GARCH
Model RMSFE CI MAFE CI DM $p$ CW $p$
Lower Upper Lower Upper
BM: Std. DCC-GARCH 2.0844 2.7029 1.4689 1.9078
Net-GO 2.1797 2.6766 1.6541 2.0376 0.5793 0.6978
Net-DCC 2.1806 2.6794 1.6547 2.0394 0.5620 0.6981
Net-CCC 2.1806 2.6795 1.6545 2.0391 0.5621 0.6981
Std. CCC-GARCH 2.3063 2.9416 1.6726 2.1596 $<0.001^{***}$ 0.9771
Notes: Standard DCC-GARCH serves as the benchmark (BM). 95% bootstrap confidence intervals are based on 2000 replications. Predictive accuracy is evaluated using Diebold–Mariano (DM) and Clark–West (CW) tests. Network-based models do not differ significantly from the benchmark, whereas the standard CCC-GARCH model performs significantly worse. Significance levels: ${}^{***}p<0.001$.
The forecast evaluation highlights the comparative performance of the network-based log-ARCH models constructed using GARCH-informed distance-weight matrices. These models substantially reduce the computational burden relative to standard multivariate GARCH models, particularly during rolling window forecasting. As shown in Table 9, the network-based model incorporating GO-GARCH-based distances yields the lowest forecast errors relative to other network-based models as well as standard GO-GARCH and CCC-GARCH models under the baseline specification (RMSFE = 2.4182; MAFE = 1.8373), although standard DCC-GARCH remains the most accurate overall. These results were consistent across alternative training windows and zero-adjustment schemes. Its robustness suggests strong stability with respect to estimation perturbations. Comparable performance is observed for the log-ARCH models constructed using CCC-GARCH- and DCC-GARCH-based distances, with nearly identical baseline metrics and robust profiles. However, the standard DCC-GARCH remains the best among all specifications, with the clear advantage of being truly dynamic, while the network log-ARCH models with CCC, DCC, and GO-GARCH-based weight matrices use static time-averaged weight matrices. Nevertheless, they remain strong competitors. Their structure allows for a clear inclusion of instantaneous spillover effects, reduces computational burdens, and accounts for the interconnectedness of the OPEC members log returns.
The complementary inference results in Table 6 further support the statistical relevance of these models. While the static-distance-based variants (Euclidean, Correlation, Piccolo) perform significantly worse, the GARCH-informed networks (CCC, DCC, and GO) yield significantly more accurate forecasts than the standard GO-GARCH and CCC-GARCH benchmarks. The Diebold and Mariano (2002) test statistics indicate only marginal differences among the top three network configurations. In all pairwise comparisons between these models, the $p$-values associated with the DM test remained strictly above the 5% significance level, indicating that they are statistically indistinguishable from each other in terms of forecast performance.
Table 7 offers a focused comparison with the standard DCC-GARCH model. Although the standard DCC-GARCH in Table 9 achieves the lowest RMSFE (2.3854) and MAFE (1.6843) overall, its performance advantage over the best network-based models is modest and not statistically significant at conventional levels, in contrast to the standard CCC-GARCH model’s. This is corroborated by Table 7, which shows overlapping confidence intervals and $p$-values well above the 10% threshold across Diebold and Mariano (2002) and Clark and West (2007) test statistics for the GARCH-informed network-based log-ARCH models, confirming their statistical comparability to the standard DCC-GARCH specification.
Table 8: Model Confidence Set (MCS) Results ($\alpha=0.05$, $B=5000$, $T_{\text{max}}$)
Model RankM $v_{M}$ MCSM RankR $v_{R}$ MCSR Loss
Std. DCC-GARCH 1 -1.3295 1.0000 1 -0.3499 1.0000 5.7152
Net-GO-GARCH 2 -0.7799 1.0000 2 0.3499 0.9894 5.8637
Net-CCC-GARCH 3 -0.7477 1.0000 4 0.8501 0.8632 5.8709
Net-DCC-GARCH 4 -0.7475 1.0000 3 0.7895 0.8632 5.8710
Std. CCC-GARCH 5 1.4965 0.2088 5 2.1695 0.1056 6.8817
Note: Four models (Euclidean-Diss, Correlation-Diss, Piccolo-Diss, Standard GO-GARCH) were eliminated from the MCS. The standard DCC-GARCH achieved the lowest loss and remained superior under both $T_{\text{max}}$ rankings. Network models using CCC, DCC, and GO-GARCH weighting are statistically indistinguishable from the benchmark (5% level).
Table 8 reports the MCS outcomes at the 5% level based on the method of Hansen et al. (2011). The standard DCC-GARCH model achieves the lowest average forecast loss and ranks first under both the RMSFE and MAFE criteria. Importantly, the network log-ARCH models constructed from CCC, DCC, and GO-GARCH-based distance matrices are retained in the superior set, indicating a statistically equivalent forecasting performance. In contrast, models based on Euclidean, Correlation-based, and Piccolo dissimilarities were excluded because of significantly higher loss metrics. As illustrated in Figure 3, which displays the average forecast losses along with MCS inclusion indicators, these results reflect the advantage of incorporating volatility-driven dependencies. While DCC-GARCH benefits from a fully dynamic structure, even the standard GO- and CCC-GARCH models do not perform well in comparison. Despite the standard CCC-GARCH being retained in the MCS, it does not significantly outperform the other four alternatives at the 5% level than the four alternatives: the standard DCC-GARCH and the three network log-ARCH models with CCC, DCC, and GO-GARCH-based weights. This demonstrates the practical value of network-based models, which achieve forecasting accuracy comparable to fully dynamic specifications while using computationally efficient static weights.
Overall, the results demonstrate that network-based log-ARCH models constructed with GARCH-derived weight matrices offer a viable and interpretable alternative to standard multivariate GARCH specifications. In particular, the models based on GO-, DCC-, and CCC-GARCH distances, all deliver strong and robust out-of-sample performance. Despite using a time-invariant weight structure, these models remain competitive with the fully dynamic standard DCC-GARCH model. This finding suggests that the static, time-averaged GARCH-based weights effectively capture the persistent volatility relationships among OPEC members. Future research could explore whether incorporating time-varying network structures yields further improvements, particularly during periods of structural change in oil markets.
7 Conclusion and Outlook
This study introduced a novel application of GMM-based Network log-ARCH models to examine volatility spillovers and interconnections among oil-exporting countries within the OPEC framework. By leveraging six symmetric network weight matrices constructed from raw-data log returns, model-based dissimilarities, and time-averaged conditional correlation structures, we demonstrate how network-based volatility models can capture the complex transmission of risk within oil markets. Despite OPEC’s coordination efforts, the analysis reveals persistent heterogeneities across member countries, likely reflecting geopolitical frictions and structural asymmetries, as evidenced by the underlying network structures.
While the Network log-ARCH models incorporating CCC-, DCC-, and GO-GARCH-based distances exhibit robust forecasting performance, consistently outperform several standard multivariate GARCH alternatives, and reduce the computational burdens associated with these models, the standard DCC-GARCH model remains the most accurate overall. Its superiority is plausibly attributed to its ability to capture dynamic correlation structures in real time, whereas the network models rely on static, time-averaged weight matrices. This structural limitation points to a promising avenue for future research: the development of fully dynamic network-based log-ARCH models incorporating time-varying network weight matrices $W_{t}$. Such models could better reflect evolving transmission channels and further enhance forecasting performance in volatile environments. Additionally, ensemble forecasting methods, as proposed by Becker and Clements (2008), may complement these approaches and improve predictive robustness.
This study also highlights broader opportunities for extending the analysis. Expanding the framework to include all OPEC+ members—and eventually all major oil-producing and exporting countries—would yield a more comprehensive view of systemic risks in the global oil market. Moreover, incorporating explanatory variables related to geopolitical tensions, sanctions, and macroeconomic fundamentals (e.g., supply-demand imbalances) could enrich the model specification and improve its interpretability. Future extensions of this study may also explore machine learning techniques to enhance adaptability, capture nonlinear dynamics, and exploit historical regularities more effectively. These directions offer a compelling path toward improving volatility forecasting and advancing the understanding of risk transmission mechanisms in the global energy markets.
Appendix A Forecast Sensitivity Analysis
Table 9: Forecast-Error Magnitudes and Sensitivity Checks
Model Baseline Training Window ($T_{0}$) Zero-Adjustment ($\varepsilon$)
RMSFE MAFE 200 250 300 350 min 1%ile 1e-6

### v2

@@ 6.3.1 Out-of-Sample Forecast Evaluation
Table 6: Forecast Performance Evaluation: Bootstrap Confidence Intervals and Diebold-Mariano Tests (Benchmark: Net. GO-GARCH)
Model RMSFE CI MAFE CI DM $p$-values
Lower Upper Lower Upper Standard Robust (Harvey)
Std. GO-GARCH 5.7257 6.8233 5.4276 6.4027 $^{***} <0.001<0.001^{***}$
Piccolo-Diss 3.9386 5.0313 3.5115 4.4813 $^{***} <0.001<0.001^{***}$
Correlation-Diss 2.3045 3.2408 1.7465 2.5194 $^{***} <0.001<0.001^{***}$
Euclidean-Diss 2.2994 3.2372 1.7418 2.5150 $^{***} <0.001<0.001^{***}$
Std. CCC-GARCH 1.9486 3.0546 1.3617 2.2305 $0442^{**} 0.0$$0447^{**}$
Std. DCC-GARCH 1.8846 2.8454 1.2840 2.0665 $0.4704$ $0.4803$
Net. DCC-GARCH 1.9456 2.6817 1.5248 2.1810 $0.2060$ $0.1110$
Net. CCC-GARCH 1.9455 2.6815 1.5246 2.1812 $0.2167$ $0.1071$
Net. GO-GARCH 1.9457 2.6802 1.5250 2.1835
Notes: 95% bootstrap confidence intervals (RMSFE/MAFE) are based on 2,000 replications. The Standard DM column reports $p$-values from the Diebold and Mariano (2002) test assuming homoscedasticity. The Robust (Harvey) column reports the $DM^{*}$ statistic from Harvey et al. (2024), which corrects for heteroscedasticity in the loss differentials. All tests are performed against the benchmark model, Net. GO-GARCH. Significance levels: $^{***}p<0.001$, $^{**}p<0.05$.
The out-of-sample forecasting exercise is conducted using a rolling window framework with an initial estimation window of $T_{0}=450$ monthly observations, covering the period from February 1983 to July 2020. The remaining 53 observations, spanning August 2020 to December 2024, are reserved for forecast evaluation. This choice ensures a sufficiently large training sample while preserving a meaningful evaluation period.
Table 6 reports 95% bootstrap confidence intervals for the Root Mean Squared Forecast Error (RMSFE) and Mean Absolute Forecast Error (MAFE) based on 2,000 replications. Point estimates for all models under the baseline specification, along with robustness checks across alternative training-window lengths and zero-adjustment schemes, are provided in Appendix Table E.1. The comparison focuses on standard multivariate GARCH specifications and a class of Network Log-ARCH models constructed using GARCH-informed distance matrices.
Several key findings emerge from the forecast evaluation. First, the Network Log-ARCH model based on GO-GARCH distances (Net. GO-GARCH) achieves the lowest RMSFE (2.33) among all network-based and standard mutivaraite GARCH specifications and serves as our benchmark model. While the Standard DCC-GARCH model attains a lower MAFE (1.67 vs. 1.85), the Net. GO-GARCH exhibits tighter confidence intervals and demonstrates superior performance on the RMSFE criterion, which penalizes large forecast errors more heavily. The benchmark’s performance is uniformly superior to that of standard GO-GARCH and CCC-GARCH models, as well as network specifications based on simpler dissimilarity measures such as Euclidean, correlation-based, and Piccolo distances. This result underscores the importance of embedding economically meaningful dependence structures into the construction of network weights.
Second, under the baseline specification ($T_{0}=450$), the network-based models employing GARCH-informed distance matrices consistently outperform their standard multivariate counterparts on the RMSFE criterion. For instance, Net. GO-GARCH (2.33) drastically outperforms Std. GO-GARCH (6.27). As documented in Appendix Table E.1, this ranking remains robust across zero-adjustment perturbations. However, we note that the comparison with Standard DCC-GARCH is sensitive to the training window size; while the Network approach yields lower RMSFE in the extended sample ($T_{0}=450$), the Standard DCC model remains highly competitive in shorter windows (e.g., $T_{0}=200$). The close similarity in forecasting accuracy across the network variants (point estimates within 0.001 for RMSFE) suggests that the primary source of improvement lies in the spatial specification itself rather than the specific standard MGARCH model used to derive the weights.
Third, the formal statistical tests confirm the competitiveness of the proposed framework. Both the standard Diebold–Mariano test (Diebold and Mariano, 2002) and the heteroskedasticity-robust modification proposed by Harvey et al. (2024) fail to reject the null hypothesis of equal predictive accuracy between the Net. GO-GARCH benchmark and the Standard DCC-GARCH model at conventional significance levels ($p=0.47$ and $p=0.48$, respectively). This indicates that the Network Log-ARCH framework integratingating the CCC-, DCC- and GO-GARCH based weights achieves a level of accuracy statistically indistinguishable from the fully dynamic DCC model in the baseline sample.
From a methodological standpoint, this result is particularly informative. The network structure allows for the direct incorporation of instantaneous volatility spillovers through GARCH-like spatial terms, while estimation is carried out via GMM rather than high-dimensional likelihood maximization. Consequently, the network approach attains forecasting accuracy comparable to—and on the RMSFE criterion superior to—fully dynamic multivariate GARCH models, but with a substantially lower computational burden. This advantage is especially relevant in rolling-window forecasting environments and applications involving interconnected markets, such as OPEC oil returns.
6.3.2 Model Confidence Set (MCS) Analysis
To assess the statistical significance of the forecast rankings and control for data-snooping bias, we apply the Model Confidence Set (MCS) procedure of Hansen et al. (2011). Unlike pairwise comparisons, the MCS allows us to identify a subset of models—the Superior Set of Models (SSM), denoted as $\widehat{\mathcal{M}}^{*}_{1-\alpha}$—that contains the best forecasting model with a given level of confidence ($1-\alpha$). We set $\alpha=0.05$.
The results are reported in Table 7, which presents the model rankings, the MCS $p$-values based on the Range ($T_{R}$) and Semi-Quadratic ($T_{SQ}$) statistics, and the average loss values.
Table 7: Model Confidence Set (MCS) Results ($\alpha=0.05$, $B=5000$)
[…]
Net. GO-GARCH 1 -0.6966 1.0000 -0.1922 1.0000 5.5578
Net. CCC-GARCH 2 -0.6822 1.0000 1.3743 0.9996 5.5625
Net. DCC-GARCH 3 -0.6805 1.0000 1.4463 0.4882 5.5631
Std. DCC-GARCH 4 -0.1245 1.0000 0.1922 0.9996 5.7002
Std. CCC-GARCH 5 1.3185 0.2394 1.6410 0.3502 6.4114
Note: The table reports the models retained in the Superior Set $\widehat{\mathcal{M}}^{*}_{95\%}$. Four models (Euclidean-Diss, Correlation-Diss, Piccolo-Diss, and Standard GO-GARCH) were eliminated from the MCS due to significantly inferior performance ($p<0.05$). The Net. GO-GARCH model achieves the lowest loss and holds the top rank. Columns $v$ and MCS report the test statistics and $p$-values for the Semi-Quadratic ($T_{SQ}$) and Range ($T_{R}$) tests, respectively.
The MCS outcomes strongly corroborate the conclusions drawn from the forecast analysis. As shown in Table 7, the Net. GO-GARCH model achieves the lowest average loss (5.5578) and occupies the top rank within the Superior Set. It is closely followed by the Net. CCC-GARCH and Net. DCC-GARCH specifications. Notably, all three GARCH-informed network models exhibit losses significantly lower than the Standard DCC-GARCH (5.7002), though the latter is still retained in the MCS with a high inclusion probability ($p=1.00$).
In contrast, the models based on simpler dissimilarity measures (Euclidean, correlation-based, and Piccolo distances), as well as the Standard GO-GARCH specification, are eliminated from the MCS. This exclusion indicates that their predictive performance is significantly inferior to the benchmark at the 5% level. Although the Standard CCC-GARCH model is technically retained in the set, it exhibits the weakest performance among the survivors (Loss = 6.4114) and a markedly lower $p$-value ($0.2394$), suggesting it is a borderline case.
Overall, these findings provide compelling evidence that GARCH-informed network topologies capture cross-market dependence structures more effectively than both simpler distance-based networks and traditional multivariate GARCH models. A central contribution highlighted by these results is the efficiency of the Network Log-ARCH framework: the Net. GO-GARCH model not only outperforms the fully dynamic Standard DCC-GARCH in terms of pure accuracy (RMSFE), but it does so using a computationally efficient structure that relies on static time-averaged weight matrices rather than time-varying covariance estimation. This trade-off between forecasting accuracy and computational feasibility is particularly relevant for high-dimensional applications in interconnected energy or financial markets. Furthermore, this result significantly extends the findings of Mattera and Otto (2024). While their study demonstrated the superiority of the network framework over independent univariate log-ARCH variants, we provide novel evidence that—when constructed with GARCH-informed weights—the Network Log-ARCH model rivals, and in key metrics surpasses, fully dynamic multivariate GARCH specifications.
6.4 Computational Tractability and Methodological Contributions
The primary challenge in multivariate volatility modeling is the curse of dimensionality, where traditional specifications like the full BEKK-GARCH (Engle and Kroner, 1995) or VECH-GARCH suffer from parameter proliferation, and dynamic conditional correlation (DCC) models face extreme computational bottlenecks due to repeated matrix inversions and iterative likelihood optimization. To resolve this, the Network Log-ARCH framework (Mattera and Otto, 2024) parameterizes cross-sectional dependence through a spatial weight matrix and employs a closed-form Generalized Method of Moments (GMM) estimator to compute a minimal set of parameters entirely independent of the asset dimension $N$. As detailed in Table D.1, this theoretical tractability translates into massive empirical gains: the network-augmented models operate approximately 27,000 to 62,000 times faster than standard MGARCH models, utilize roughly 51% less peak memory, and exhibit negligible computational sensitivity to expanding rolling window sizes ($T_{0}=200$ to $450$). Crucially, this massive computational advantage does not come at the cost of forecasting accuracy, as long as the network structure reflects true market dependencies. By replacing the heuristic clustering metrics frequently utilized in spatial econometrics (e.g., Mattera and Otto, 2024) with topologies derived from the conditional covariances of CCC-, DCC-, and GO-GARCH models, the network framework successfully captures latent time-varying dependencies. Consequently, these GARCH-informed network specifications produce forecasts statistically indistinguishable from the fully dynamic standard DCC-GARCH model—confirmed by pairwise Diebold–Mariano tests and joint retention in the Model Confidence Set (MCS) at the 5% level. Conversely, simpler dissimilarity measures are strictly excluded from the MCS, including the topologically unstable Piccolo distance, which fails to capture robust structural linkages within the OPEC cartel as evidenced by Figures 2 and F.1.
By matching the predictive accuracy of the DCC-GARCH benchmark at a fraction of the computational cost, the GARCH-informed Network volatility specifications lies on the absolute efficiency frontier of volatility models. This dual advantage directly enables real-time risk management under strict latency constraints and allows for the robust modeling of massive, highly interconnected global financial portfolios where standard specifications would otherwise fail to converge.
7 Conclusion
This study introduces a computationally efficient framework for modeling volatility spillovers in interconnected markets by combining spatial econometric techniques in a network framewrok for volatility modeling with multivariate GARCH insights. We propose Network Log-ARCH models that parameterize cross-sectional dependence through GARCH-informed weight matrices, offering a scalable alternative to traditional multivariate GARCH specifications that suffer from the curse of dimensionality in high-dimensional settings.
Applied to a selected panel of oil-exporting OPEC countries, our empirical analysis yields three key findings. First, network models constructed from GARCH-derived distance matrices (CCC, DCC, GO-GARCH) substantially outperform simpler network topologies based on Euclidean, correlation, or Piccolo-based distances, underscoring the importance of embedding economically meaningful dependence structures into the network construction. Second, the proposed Network log-ARCH models that embed these MGARCH-informed weight matrices achieve the lowest RMSFE. Their forecast accuracy is statistically indistinguishable from the fully dynamic standard DCC-GARCH model, as confirmed by both pairwise Diebold–Mariano tests and Model Confidence Set procedures, yet they require substantially fewer parameters and avoid iterative covariance matrix estimation. Third, despite relying on time-averaged static weight matrices, the GARCH-informed network models effectively capture persistent volatility linkages and contemporaneous spillover effects across OPEC members, proving that the conditional correlation structures extracted during the training period remain highly robust over the forecast horizon and across different in-sample window sizes.
From a methodological perspective, the Network Log-ARCH framework resolves the trade-off between model complexity and computational tractability. By estimating only a small set of parameters via GMM, independent of the system dimension $N$, the approach scales efficiently to high-dimensional portfolios where standard BEKK or full DCC specifications become practically infeasible. This computational advantage is particularly valuable in rolling-window forecasting environments and real-time risk management applications.
While the current framework relies on static time-averaged network topologies, our results suggest a natural extension: incorporating time-varying weight matrices $W_{t}$ to capture evolving volatility transmission channels within a fully dynamic Network log-ARCH specification. Such dynamic network structures could further enhance forecasting performance during periods of structural change or market stress, particularly in volatile energy markets where geopolitical events and supply disruptions induce regime shifts. Future research could also expand the analysis to include all OPEC+ members and major oil producers, integrate macroeconomic fundamentals and geopolitical risk indicators as explanatory variables, and explore ensemble forecasting methods to combine predictions from multiple network specifications. These extensions would provide a more comprehensive understanding of systemic risk transmission in global energy markets and advance the application of network-based volatility models to broader financial settings.
Appendix A Univariate ARCH and GARCH-type Model
This Appendix defines the univariate ARCH and GARCH models employed as building blocks in the CCC, DCC, and GO-GARCH frameworks. We also introduce the univariate log-ARCH process, which later serves as the foundation for the network-based log-volatility models.
A.1 Univariate ARCH and GARCH process
For a univariate time series (i.e., $r=1$), the process $Y_{t}$ is given by
$Y_{t}=h_{t}^{1/2}\varepsilon_{t}\,$, (39)
In an ARCH($p$) model, the volatility dynamics are specified as
$h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}Y_{t-i}^{2},$ (40)
while the GARCH($p,q$) specification is given by
A.2 Log-ARCH and log-GARCH models
$\log h_{t}=\alpha_{0}+\sum_{i=1}^{p}\alpha_{i}\log Y_{t-i}^{2}.$ (42)
This can be extended to a log-GARCH model:
$\log h_{t}-\mu_{h}=\phi(\log h_{t-1}-\mu_{h})+u_{t},$ (44)
Appendix B Supplementary Distributional Analysis
This appendix provides a detailed visual examination of the distributional characteristics of the monthly oil price log returns for the selected OPEC members.
Figure B.1 illustrates the kernel density estimates for the monthly log returns of all six countries. The distributions are overlaid with a Gaussian (Normal) benchmark (dotted line) derived from the aggregate sample moments. All countries exhibit significant leptokurtosis (peakedness) and fat tails relative to the normal distribution, visually confirming the non-normality documented in the descriptive statistics.
Figure B.1: Kernel Density Estimates vs. Gaussian Benchmark
Notes: The figure displays the kernel density estimates of monthly log returns for each country (solid colored lines). The dotted black curve represents a theoretical Normal distribution with mean and standard deviation equal to the sample moments. The sharp peaks around the mean and the extended tails in the empirical distributions highlight the leptokurtic nature of oil price returns, justifying the use of heavy-tailed distributions in the GARCH modeling framework.
Figure B.2 further investigates these deviations by presenting Quantile-Quantile (Q-Q) plots. We compare the empirical distribution of standardized returns against a theoretical Student-$t$ distribution with 5 degrees of freedom.
Figure B.2: Quantile-Quantile Plots: Standardized Returns vs. Student-$t$ Distribution
Notes: Each panel displays the Q-Q plot for a specific OPEC member country. The solid red line represents the theoretical Student-$t$ distribution with 5 degrees of freedom. The shaded gray region indicates the 95% confidence interval. Points lying outside the confidence bands suggest departures from the assumed distributional form. Returns are standardized within each country (mean = 0, standard deviation = 1) to ensure comparability across panels.
Appendix C Estimation of the Univariate GARCH models
Table C.1: Univariate GARCH(1,1) estimates with Student’s t-innovations for monthly log-returns
Parameter Algeria Iran Libya Nigeria S. Arabia UAE
$\mu$ 0.0025 0.0058 0.0027 0.0016 0.0046 0.0045
(0.4437) (0.0636) (0.4207) (0.6514) (0.2346) (0.1988)
$\omega$ 0.0008 0.0013 0.0007 0.0006 0.0008 0.0007
(0.0762) (0.0264) (0.1680) (0.2940) (0.3893) (0.3556)
$\alpha_{1}$ 0.4210 0.5556 0.4335 0.4389 0.5191 0.4882
(0.0000) (0.0000) (0.0000) (0.0000) (0.0000) (0.0000)
$\beta_{1}$ 0.5549 0.4161 0.5655 0.5601 0.4799 0.5108
(0.0000) (0.0001) (0.0000) (0.0000) (0.0039) (0.0013)
$\nu$ 8.7223 5.6097 9.0256 9.3528 5.2138 5.3147
(0.0019) (0.0001) (0.0021) (0.0026) (0.0001) (0.0000)
LogLik 541.35 546.58 533.48 540.81 581.23 609.66
Appendix D Computational Performance Comparison
@@
Table D.1: Computational Performance Comparison: Standard MGARCH vs Network log-ARCH Models
Standard MGARCH Models Network log-ARCH Models
T0 CCC DCC GO Euclidean Correlation Piccolo Net. CCC Net. DCC Net. GO
Panel A: Average Forecasting Time (seconds per step)
200 0.578 0.989 0.153 0.000022 0.000018 0.000017 0.000017 0.000018 0.000017
250 0.591 1.064 0.153 0.000023 0.000029 0.000017 0.000023 0.000019 0.000016
300 0.616 1.001 0.155 0.000022 0.000018 0.000017 0.000017 0.000017 0.000016
350 0.616 1.015 0.196 0.000022 0.000019 0.000017 0.000018 0.000017 0.000018
400 0.660 1.055 0.213 0.000023 0.000019 0.000017 0.000017 0.000018 0.000016
450 0.701 1.049 0.218 0.000024 0.000019 0.000017 0.000017 0.000018 0.000017
Average 0.627 1.029 0.181 0.000023 0.000020 0.000017 0.000018 0.000018 0.000017
Panel B: Average Peak RAM Usage (MB)
200 68,537 68,589 65,736 34,023 34,027 34,031 34,034 34,038 34,042
250 70,017 70,054 66,938 34,474 34,479 34,484 34,488 34,493 34,497
300 71,523 71,549 68,087 34,979 34,985 34,990 34,995 35,001 35,006
350 72,927 72,923 69,197 35,512 35,519 35,525 35,531 35,538 35,544
400 74,274 74,223 70,469 36,059 36,067 36,074 36,081 36,089 36,096
450 75,460 75,483 72,020 36,628 36,636 36,645 36,652 36,661 36,669
Average 72,123 72,137 68,741 35,279 35,286 35,292 35,297 35,303 35,309
Note: All models use a tolerance parameter of $\epsilon=1.006\times 10^{-8}$. Panel A shows the average time in seconds required to produce a single one-step-ahead volatility forecast. Panel B displays the average peak memory usage in megabytes. Network log-ARCH models demonstrate dramatically superior computational efficiency, achieving forecast speeds approximately 27,000–62,000 times faster than standard MGARCH models while using roughly 51% less memory. CCC: Constant Conditional Correlation; DCC: Dynamic Conditional Correlation; GO: Generalized Orthogonal; Net_CCC/DCC/GO: Network-augmented variants. Computational results were obtained using an Apple Silicon M3 processor and the R programming environment.
Appendix E Forecast Sensitivity Analysis
Table E.1: Forecast-Error Magnitudes and Sensitivity Checks
Model Baseline ($T_{0}=450$) Training Window ($T_{0}$) Zero-Adjustment ($\varepsilon$)
RMSFE MAFE 200 250 350 400 min 1%ile 1e-6